Solve : y^2 + x^2 (dy/dx)^2 -2xy (dy/dx) = 4 (dx/dy)^2
1
Expert's answer
2020-02-23T16:06:53-0500

"y^2-2xyy'+x^2(y')^2-\frac{4}{(y')^2}=0\\
a=1, b=-2xy', c=x^2(y')^2-\frac{4}{(y')^2}\\
D=b^2-4ac=4x^2(y')^2-4(x^2(y')^2-\frac{4}{(y')^2})=\\
=\frac{16}{(y')^2}\\
1) y=\frac{2xy'-\frac{4}{y'}}{2} \implies y=xy'-\frac{2}{y'} (*)"

this is the Clairaut's equation, make a substitution

"y'=p, dy=p dx"

then

"y=xp-\frac{2}{p}\\
dy=pdx +xdp+\frac{2}{p^2}dp\\
pdx=pdx +xdp+\frac{2}{p^2}dp\\
(x+\frac{2}{p^2})dp=0\\
a) dp=0 \implies p=c \implies y=xc-\frac{2}{c}"

is the general solution

"b) x+\frac{2}{p^2}=0"

let's say "x<0" then

"p=\pm\sqrt\frac{2}{-x} \implies y'=\pm\sqrt\frac{2}{-x} \implies\\
y=\pm(-2\sqrt2 \sqrt{(-x)}) +c_1"

substitute y in (*)

"\pm(-2\sqrt2 \sqrt {( -x )})+c_1=\\
=x(\pm\sqrt\frac{2}{-x})-(\pm\frac{2}{\sqrt\frac{2}{-x}})\implies\\
c_1=0"

then "y=\pm(-2\sqrt2 \sqrt{(-x)})" are solutions of the equation.

2) "y=\frac{2xy'+\frac{4}{y'}}{2} \implies y=xy'+\frac{2}{y'} (**)"

this is the Clairaut's equation, make a substitution

"y'=p, dy=p dx"

then

"y=xp+\frac{2}{p}\\
dy=pdx +xdp-\frac{2}{p^2}dp\\
pdx=pdx +xdp-\frac{2}{p^2}dp\\
(x-\frac{2}{p^2})dp=0\\
a) dp=0 \implies p=c \implies y=xc+\frac{2}{c}"

is the general solution

"b) x-\frac{2}{p^2}=0"

let's say "x>0" then

"p=\pm\sqrt\frac{2}{x} \implies y'=\pm\sqrt\frac{2}{x} \implies\\
y=\pm2\sqrt2 \sqrt x +c_2"

substitute y in (**)

"\pm2\sqrt2 \sqrt x+c_2=\pm x\sqrt\frac{2}{x}+\frac{2}{\pm\sqrt\frac{2}{x}}\implies\\
c_2=0"

then "y=\pm2\sqrt2 \sqrt x" are solutions of the equation.